OpenStudy (anonymous):

A manufacturer produces bolts of a fabric with a fixed width. A quantity q of this fabric (measured in yards) that is sold is a function of the selling price p (in dollars per yard), so we can write q=f(p). Then, the total revenue earned with selling price p is R(p)=pf(p). Find R′(20), given f(20)=10000, and f′(20)=−350. (What does this mean?)

OpenStudy (anonymous):

R'(p) = (pf(p))' = p' * f(p) + p * f'(p) = f(p) + p*f'(p) plug p = 20 and get R'(20) = f(20) + 20 * f'(20)

OpenStudy (anonymous):

can you do this now ?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

yw